Circular No. 8416

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2004 T1 (LINEAR-NEAT) K. J. Lawrence, Jet Propulsion Laboratory, reports the NEAT discov-

ery of a comet on images taken with the Haleakala telescope, the object described as having a coma of diameter ∼ 25′′, with elongated images but no obvious tail. The comet was also observed earlier on the same night (with prediscovery observations on Sept. 21) by the LINEAR team, which made no comment on the object’s appearance; both “discovery” observations are tabulated below. Following posting on “The NEO Confirmation Page”, numerous observers have reported on the object’s cometary appearance on CCD images obtained during Oct. 5.9–6.6 UT, including M. Tichý (Kleť), P. Birtwhistle (Great Shefford, U.K.), C. Jacques and E. Pimentel (Belo Horizonte, Brazil), D. T. Durig and G. A. T. Morris (Sewanee, TN), J. E. McGaha (Tucson, AZ), G. R. Jones (Tucson, AZ), and D. Higgins (Canberra, Australia). The general consensus is that there is a diffuse coma of diameter ∼ 16′′–30′′ and a faint, broad, fan-shaped tail ∼ 1′ wide and 4′–5′ long in p.a. ≈ 240^{o}–280^{o}. P. Kušnirák (Ondřejov) reports the fan tail in p.a. ∼ 290^{o} and another tail ∼ 4′ long in p.a. ∼ 240^{o}.

2004 UT α_{2000} δ_{2000} Mag. Observer
Oct. 5.26504 0^{h}57^{m}02.^{s}82 −12^{o}15′12′′.0 17.2 LINEAR
5.43686 0 56 55.35 −12 14 01.1 15.2 NEAT

The available astrometry, the following preliminary orbital elements, and an ephemeris appear on MPEC 2004-T18 (H_{10} = 13.0).

T = 2004 Nov. 5.517 TT  ω = 334.^{o}935
e = 0.50537  Ω = 51.733 }2000.0
q = 1.71781 AU  i = 11.013
a = 3.47290 AU n^{o} = 0.152288 P = 6.47 years
2004TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  Mag.
Oct. 2 0^{h}59.^{m}28 −12^{o}34.^{′}7 0.767 1.748 162.^{o}6 9.^{o}9 14.8
12 0 52.31 −11 23.1 0.759 1.733 160.7 11.0 14.8
22 0 46.00 − 9 36.9 0.772 1.723 154.3 14.5 14.8
Nov. 1 0 41.91 − 7 20.1 0.802 1.718 146.3 18.7 14.9
11 0 41.00 − 4 41.0 0.850 1.719 138.0 22.7 15.0
21 0 43.65 − 1 47.9 0.913 1.724 130.1 26.0 15.2
Dec. 1 0 49.77 + 1 11.7 0.990 1.734 122.8 28.6 15.4
2004 October 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green